relation btn Linear and volumetric thermal coefficient 

Grade:11

1 Answers

Arun
5 years ago
Dear Atul
 

α = coefficient of linear expansion

β = coefficient of areal expansion

γ = coefficient of volume expansion

Since, coefficient of linear expansion is in one direction and areal in two dimension followed by three dimension in case of volume expansion. So, to compensate that factor the relation between them will be like.

α = β /2= γ /3